The ISS in the orbit 408 km away from the surface of the earth. Calculate the time taken to orbit the earth. 



1
Expert's answer
2021-05-27T17:47:55-0400
"T=2\pi\sqrt{\frac{R^3}{GM}}\\T=2\pi\sqrt{\frac{(408\cdot10^{3}+6371\cdot10^{3})^3}{(6.67\cdot10^{-11})(5.97\cdot10^{24})}}=5560\ s"